OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

so to get from the 2nd term to the 5th term we have to multiply by the same number "r", 3 times once to get the 3rd term again to get the 4th term and lastly to get the 5th term so \(\begin{array}{llll} term&value \\\hline\\ a_2&36\\ a_3&36\cdot r\\ a_4&(36\cdot r)\cdot r\\ a_5&(36\cdot r\cdot r)\cdot r\to 2,304 \end{array}\) so we can say that \(\bf 36r\cdot r\cdot r=2,304\implies 36r^3=2,304\implies r=\cfrac{2,304}{36}\)

OpenStudy (jdoe0001):

hmm actually I missed the 3 =) heheh lemme fix that

OpenStudy (jdoe0001):

\(\bf 36r\cdot r\cdot r=2,304\implies 36r^3=2,304\implies r^3=\cfrac{2,304}{36} \\ \quad \\ \implies\large r=\sqrt[3]{\cfrac{2,304}{36}}\)

OpenStudy (anonymous):

is that complete?

OpenStudy (jdoe0001):

well... no... .but what would that give us for "r"?

OpenStudy (anonymous):

oh, let me solve.

OpenStudy (anonymous):

r=4

OpenStudy (anonymous):

so my answer for this problem would be an = 9 • 4^(n - 1) ?

OpenStudy (jdoe0001):

yeap so if r = 4 and the 2nd term is 36 that means that 4 * "something" is 36 and that "something" is our first term so our first term will be \(\bf "something"\cdot 4=36\implies "something"=\cfrac{\cancel{ 36 }}{\cancel{ 4 }}\to 9\leftarrow a_1\) so now we know our \(\bf a_1\ and \ r\) so just plug them in :) \(\Large a_{\color{brown}{ n}}=a_1\cdot r^{{\color{brown}{ n}}-1}\)